The base bias circuit is subjected to a temperature variation from 0°C to 70°C. The ß decreases by 50% at 0°C and increases by 75% at 70°C from it's normal value of 110 at 25°C. What are the changes in IC and VCE over the temperature range of 0°C to 70°C Given: RB = 15kΩ, Rc = 100Ω, Vcc = 9V
1
Expert's answer
2020-10-05T15:17:55-0400

Initialy we have base bias circuit with given quantities as


"R_B=15k\Omega,R_c=100\Omega,V_{cc}=9V" ,"\beta" at 25"\degree C" = 110


Since the given circuit is a base bias circuit then

"V_{cc}=I_B\times R_B"


"9=I_B\times 15"

"I_B=0.6mA"


Case 1: At 0"\degree C"

"\beta" reduce by 50%

so It will became "\frac{\beta}{2}"


Using formula of current gain we have,

"\frac{\beta}{2}=\frac{I_c}{I_B}"


"I_C=I_B\times \frac{110}{2}"

="0.6\times 55"

=33mA=0.033A

Now using the equation

"V_{CE}=V_{CC}-I_CR_C"

="9-0.033\times 100"

=9-3.3

=5.7V

Case 2: At "25\degree C"

"\beta=\frac{I_c}{I_B}"


"I_c=110\times0.6"

=66mA=0.066A


Again using the equation,

"V_{CE}=V_{CC}-I_CR_C"

= "9-.066\times100"

=9-6.6

=2.7V

Case 3: At "70\degree C"


"\beta" increase by 70%

i.e. "\beta" will became "\beta+\frac{75}{100}\beta"

i.e."\frac{7}{4}\beta"

using

"\frac{7}{4}\beta=\frac{I_c}{I_B}"


"I_c=I_B\times \frac{7}{4}\times 110"

"=0.6\times 7\times27.5"


=115.5ma=0.1155A

Again using,

"V_{CE}=V_{CC}-I_CR_C"

="9-0.1155\times100"

=9-11.55

"=-2.55V"